please help! thank you !
Fill in the chort and sketch the graph: rocoso-2 e cose cose ? 2400 270° 300 330° 7600
Answer #1

Caso o = C302를 명= -2 1-2 = S - 2 올 터를 --2 =-2 6일 990 = 0 에0일 (150 = -√3 에 0 터 6210 = -√3 G240 역 ... (990 = o 4300= 는 CG330 by